# Learning to combine top-down context and feed-forward representations under ambiguity with apical and basal dendrites

### v2

@@ 2.4 Deriving context from temporal information
Contextual priors can also be extracted from the temporal domain, specifically by leveraging past information to decode incoming ambiguous sensory inputs. To extend our model to cases where contextual signals are derived from the temporal domain, we trained a Gated Recurrent Unit (GRU) network to predict the arithmetic sum of two unambiguous digits as encoded by our pretrained feedforward weights $\mathbf{\Theta_{b}}$. The final output state of the GRU was then given as a contextual signal for our model. (fig. 5a; see Methods). We find that in cases where the input is ambiguous and the temporal sequence sums to a plausible interpretation of the ambiguous image, the model updates perceptual representations using contextual signals to successfully resolve ambiguities (fig. 5b, top). Importantly, context is effectively ignored in cases where input is unambiguous and contextual signals irrelevant (fig. 5b, bottom). Using low-dimensional projections of latent representations, we find that top-down context derived from the temporal domain effectively disentangled overlapping representations when inputs are ambiguous, similarly to the oracle case. (fig. 5c)
@@ 3 Discussion
The ability to appropriately use and switch contexts to make sense of the vast stream of sensory information is an important feature of robust cognitive systems. It is thought that in the neocortex, contextual representations in higher order regions modulate lower sensory regions through top-down interactions arriving on the apical dendrites of pyramidal neurons. While it is known that apical dendrites modulate neuronal firing, their specific computational role in cognitive processes such as contextual integration remains poorly understood.
Here, we tested whether the apical compartment represents an architectural bias that is functionally useful for contextual integration in neural networks. To this end, we implemented (1) a neural network architecture with distinct apical compartments and a neuron-wise integration rule based on the observation that apical dendrites are gain modulators of somatic activity (Waters et al., 2003; Larkum et al., 2007), and developed (2) an ambiguous image classification task that requires contextual integration to be solved. We trained (1) the contextual modulator (top-down synapses onto apical dendrites) by gradient descent to simultaneously learn all input-context conditions defined in our task. We found that this bio-inspired dendritic architecture outperformed a single-compartment model which processes both sensory and contextual information in the compartment. In this task setting, we find there is an inherent benefit to have contextual information propagated via a distinct apical compartment.
By applying LRP, we find that the learned top-down modulation of pyramidal neurons is consistent with the principles of biased competition (Spratling, 2008). Specifically, by analyzing the apical amplitude on class-specific neurons for ambiguous stimuli, we found the highest gain modulation in class-specific neurons that matched the target class. Moreover, we found that the level of competition introduced by the context adapted to the level of ambiguity in the task: The top-down network learned to a) refine ambiguous representations by applying strong but sparse gain modulation to basal signals when contextual information is both available and relevant, and b) provide weak modulation in the case of contradictory or irrelevant contexts or unambiguous stimuli.
@@ 3.1 Comparison with existing computational frameworks
From a computational perspective, apical dendrites have been previously shown to support some types of contextual integration in biophysically realistic neurons and ANNs (Naumann et al., 2022; Wybo et al., 2022; George et al., 2020; Baronig and Legenstein, 2024). The tasks used in these previous works are multi-task, single scenario problem settings, meaning that context (e.g. one-hot encoded task id) is always assumed to be relevant. Alternative approaches to flexible contextual integration have involved explicitly learning to prevent associations between irrelevant context and sensory inputs ((Baronig and Legenstein, 2024)), which represents a successful example of how to handle spurious contexts in a simple task. As an extension of this approach, our model must infer how to handle the interaction between sensory input and context across distinct tasks and scenarios, where context is relevant, irrelevant, or contradictory.
@@
Crucially, we only found a significant advantage of the apical prior over other integration approaches in the scenario where inputs are unambiguous and context is contradictory (and should be ignored), with an accuracy increase of 1.9% (MNIST) for models with distinct apical compartments. This highlights a specific advantage for apical dendrites in the role of contextual integration, where contextual relevance is crucial to performance. Furthermore, while the task can be solved using either the Hadamard or additive rule (Appendix 4), we found that, under uncertainty, the Hadamard integration rule enabled the model to ignore contradictory contextual signals more effectively compared to additive integration.
@@
Another key difference with most previous work is that as basal and apical signals are combined post-activation, our rule can only amplify already active neurons, unlike Wybo et al. (2022) and (Naumann et al., 2022), but inline with experimental evidence (Larkum et al., 1999). From an efficiency perspective, this maintains the sparsity of the network.
@@
One notable prior work which applies context-dependent modulation to feedforward networks is context-dependent gating (XdG) (Masse et al., 2018), which could be comparable with our basal-only ablation as it is a feedforward architecture. The main difference with our basal-only ablation is that XdG a) is not informed by intermediate sensory representations, and as most prior works, assumes context always agrees with sensory input. Therefore, its is unlikely that XdG, implemented as-is, can deal with cases where context should be overridden by sensory input, as in the unambig.Contra case.
@@
The advantage of apical dendrites are clear in our multi-scenario contextual integration task. In order to compare our results with previous work, we examined the effect of our integration mechanism in a standard multi-task setting (given task-ids). Under these conditions we found that that all variants perform similarly well. This can be explained by the setting being simpler, whereas the distinct apical compartment shines under more varied and challenging types of contextual integration that require inference of ambiguity.
While not explicitly modular, our framework could be seen as in accordance with the Global Workspace Theory along with the "information bottleneck" (to discard irrelevant information, which can include contextual information) Goyal et al. (2022). Under both frameworks, representations can be learned in a distributed and scalable manner, and mapped to local populations as context signals (in our case via apical compartments), and through biased competition, integrated with bottom-up representations to solve perceptual tasks. As highlighted in their work, this frees up the capacity of a learning system for more tasks, as it can more efficiently reuse what it has previously learned and identify task-relevant information, compared to single-compartment, single-integration models.
3.2 Predictions: What is necessary for robust contextual integration?
@@
Until now, we have shown that in our contextual integration task, the apical prior provides a specific advantage compared to single-compartment integration when sensory or contextual signals are unreliable. Here, we aim to extend these findings to make broader predictions for what specific components/computations might be necessary for robust contextual integration, and how such a mechanism could plausibly be implemented in the cortex.
@@
Our results show that, given both sensory input and contextual input must be considered simultaneously, it may be sub-optimal to sum all activity together in the basal compartment, when sensory information and context are contradictory. At a higher level, this highlights a key difference between single and two-zone integration of information. As such, an interesting experiment would be to compare the single vs two-zone integration of bottomup representations and topdown contextual representations in simulations of neurons which more faithfully respect the biophysical constraints of pyramidal neurons in the neocortex, to observe whether our results hold in a more biologically plausible setting, or if it is strictly a feature of ANNs.
@@
Furthermore, based on our results, we can infer that the apical compartment encodes some information about the ambiguity of the sensory input, and the relevance of context for a given sensory input. An interesting experiment could involve silencing or deactivating apical dendrites in more realistic neurons (biophysical simulation or biological), and observing the effect on accurate perception in a similar contextual integration task.
@@
Mutual information between apical and basal activation is positive for the ambig.Match scenario, where we see a reduction in entropy of the prediction after observing the contextual signal. On the other hand, mutual information is near 0 for the unambig.Contra scenario, which aligns with our main findings (see Appendix, Table 3). This is not surprising given that we performed supervised training. We explicitly trained under the assumption that local representations, given mis-matching context, should be invariant to sensory ambiguity. This could be interpreted as being surprised by a stimulus, but not having informative or relevant context to reduce the ambiguity of the stimulus. This is important for robust contextual integration if one subscribes to the hypothesis that, most available contextual signals in rich, real-world sensory experience are distractors for any one particular task, and should be discarded.
Following this, we predict that a model cannot learn the relevance of context with out encountering scenarios with irrelevant context during training. As an ablation study, we could remove the unambig.Irrel scenario from our loss function. We expect to observe a significant drop in accuracy on that scenario, reflecting the inability to assess contextual relevance.
@@
Extending this idea to unsupervised learning, it is interesting to think about how the model could identify contextual relevance without a supervised signal and be invariant to sensory ambiguity under unhelpful context. In an unsupervised setting this could arise from an information-maximization learning rule like InfoNCE, which uses a contrastive loss with positive and negative pairs defined via augmentation of the input (positive) against other inputs (negative). Unfortunately, the creation of positive and negative pairs relies heavily on domain-specific knowledge for useful data augmentations (eg. a rotation applied on an image). If it is possible to create contrastive pairs in a domain-agnostic manner, such an unsupervised learning framework could approximate the effects of supervised training shown in our results.
Given the fact that most associations of context and input are either irrelevant or contradictory, and only a small subset of associations represent matching context, it is not surprising that the gain modulation is sparse. We expect also that our high-dimensional apical weight matrices could be well approximated with low-rank matrices without losing significant performance, in line with work in modulating deep learning models for downstream tasks via low-rank adapters Hu et al. (2021). While this method uses resources efficiently, it lacks the mechanism to separate irrelevant or contradictory from useful contexts that the apical prior excels at, which indicates a potential synergy between the two. We hypothesize the brain uses both mechanisms to learn to handle the large pool of available sources of context efficiently. A future direction could combine LORAs with the apical prior for contextual integration tasks.
@@
We find that our model could learn a mapping of top-down signals onto the apical compartment, which importantly, was learned separately from the contextual representations. This is compatible with high-order cortical regions such as the hippocampus, where the learning and updating of contextual representations is disentangled from how they modulate sensory representations (Maren et al., 2013). Because we assume the stability of the sensory and contextual representation, we cannot make any conclusions about how these two representations could be simultaneously co-learned. However, we speculate there should be some degree of stability in the sensory representations to learn to perform contextual integration.
@@ 3.3 Limitations
While our model unifies the function of contextual integration with key biophysical properties of dendrites, some key model assumptions and simplifications are worth discussing with respect to their empirical support.
Firstly, our model is designed to implement some key features of cortical pyramidal neurons without focusing on a specific subpopulation that make up cortical layers, including layer 2/3 and layer 5 pyramidal neurons. Here, we assume that contextual signals are provided by neurons in a higher-order region, which would likely correspond to layer 5 output, and integrated in apical dendrites of putative layer 2/3 pyramidal neurons (Bastos et al., 2012; Schuman et al., 2021). It is noteworthy though that layer 5 neurons can also send top-down projections to the basal compartment of layer 2/3 neurons, although to a lesser extent (Harris et al., 2019). Unlike other computational frameworks such as (George et al., 2020) which seek to incorporate the rich architectural diversity of the neocortex, we sought to simplify and abstract many features (using neural networks) in order to be able to train models to solve more challenging contextual integration tasks which could give useful insights that could not be seen with simpler tasks.
Our model heavily abstracts temporal dynamics that are normally observed in biological circuits. For example, in the presence of ambiguous stimuli, human subjects can oscillate between two interpretations, a phenomenon termed ’perceptual bistability’ (Hardstone et al., 2021). During perceptual bistability, neural activity alternates between discrete states (Abeles et al., 1995; Akrami et al., 2009), following principles of continuous attractor dynamics (Daelli and Treves, 2010).
@@ 3.4 Outlook
In the future, our model could support more realistic temporal dynamics, which would enable the testing of a rich array of hypotheses regarding contextual integration.
In the future, we propose that our model could be extended to hierarchies of sensory modalities interacting with each other. Deep learning models that implement multi-modal interactions are an active area of research (Radford et al., 2021; Alayrac et al., 2022; Mustafa et al., 2022). While they are very good at multimodal generation with conditioning on one modality, they do not currently have the ability to resolve conflicting information and can be easily confused by contradictions in mixed-modality data. While these models employ transformer architectures that integrate context differently than MLPs, our loss function could inspire new approaches for robust multi-modal perception in deep neural networks.
From a neurophysiology perspective, our framework could be leveraged to test specific hypotheses of contextual integration in health and disease, and in particular in cases where the physiological properties of pyramidal neurons are known to be altered (Kim et al., 2014).

### v3

@@ 2.4 Deriving context from temporal information
Contextual priors can also be extracted from the temporal domain, specifically by leveraging past information to decode incoming ambiguous sensory inputs. To extend our model to cases where contextual signals are derived from the temporal domain, we trained a Gated Recurrent Unit (GRU) network to predict the arithmetic sum of two unambiguous digits as encoded by our pre-trained feedforward weights $\mathbf{\Theta_{b}}$. The final output state of the GRU was then given as a contextual signal for our contextual integration model. (fig. 5a; see Methods). We find that in cases where the input is ambiguous and the temporal sequence sums to a plausible interpretation of the ambiguous image, the model updates perceptual representations using contextual signals to successfully resolve ambiguities (fig. 5b, top). Importantly, context is effectively ignored in cases where the input is unambiguous and the contextual signals are irrelevant (fig. 5b, bottom). Using low-dimensional projections of latent representations, we find that top-down context derived from the temporal domain effectively disentangled overlapping representations when inputs are ambiguous, similar to the oracle case. (fig. 5c)
@@ 3 Discussion
The ability to appropriately use and switch contexts to make sense of a vast stream of sensory information is an important feature of robust cognitive systems.
It is thought that in the neocortex, contextual representations in higher order regions modulate lower sensory regions through top-down interactions arriving on the apical dendrites of pyramidal neurons (Larkum, 2013; Phillips, 2017), an architectual bias we refer to here as the "apical prior." While it is known that apical dendrites modulate neuronal firing, their specific computational role in cognitive processes such as contextual integration remains poorly understood.
Here, we tested whether the "apical prior" represents an architectural bias that is functionally useful for contextual integration in neural networks. To this end, we implemented (1) a neural network architecture with distinct apical compartments and a neuron-wise integration rule based on the observation that apical dendrites are gain modulators of somatic activity (Waters et al., 2003; Larkum et al., 2007), and developed (2) an ambiguous image classification task that requires contextual integration to be solved. We trained (1) the contextual modulator (top-down synapses onto apical dendrites) by gradient descent to simultaneously learn all input-context conditions defined in our task. In this setting, we found that this bio-inspired dendritic architecture outperformed a single-compartment model which processes both sensory and contextual information with a single non-linear activation.
By applying LRP, we analyzed the amplitude of apical inputs onto class-specific neurons for ambiguous stimuli. We found the highest gain modulation in context-relevant neurons (matching the target class), consistent with the principles of biased competition (Spratling, 2008).
Moreover, we found that the competition introduced by the contextual integration adapted to the level of ambiguity in the task: The top-down network learned to a) refine representations of ambiguous stimuli by applying strong but sparse gain modulation to basal signals when contextual information is both available and relevant, and b) provide weak and non-specific modulation in the case of unambiguous stimuli and contradictory or irrelevant contexts, thus preserving the bottom-up representations.
@@ 3.1 Comparison with existing computational frameworks
From a computational perspective, apical dendrites have been previously shown to support some types of contextual integration in biophysically realistic neurons and ANNs (Naumann et al., 2022; Wybo et al., 2022; George et al., 2020; Baronig and Legenstein, 2024). The tasks used in these previous works are multi-task, single scenario problem settings, meaning that context (e.g. one-hot encoded task id) is always assumed to be relevant. Alternative approaches to flexible contextual integration have involved explicitly learning to prevent associations between irrelevant context and sensory inputs (Baronig and Legenstein, 2024), which represents a successful example of how to handle spurious contexts in a simple task. As an extension of this approach, our model must infer how to handle the interaction between sensory input and context across distinct tasks and scenarios, where context is relevant, irrelevant, or contradictory, but without explicit knowledge of the task.
@@
Crucially, we only found a significant advantage of the apical prior over single compartment integration in the scenario where inputs are unambiguous and context is contradictory (and should be ignored). This highlights a specific advantage for apical dendrites in the role of contextual integration, where accounting for varying contextual relevance is crucial to broadly robust performance. Furthermore, while we show that the task can be solved using either a gain modulation (Hadamard) or additive apical integration rule (Appendix 1 Table. 4), we found that, under low contextual certainty, the Hadamard integration rule (apical prior) enabled the model to ignore contradictory contextual signals more effectively compared to additive apical integration.
@@
Another key difference of our study with most previous work is that as basal and apical signals are combined post-activation, our rule can only amplify already active neurons, unlike Wybo et al. (2022) and (Naumann et al., 2022), but inline with experimental evidence (Larkum et al., 1999). Thus, our approach maintains the sparsity of the network representations, which can preserve selectivity, and would also be more metabolically efficient.
@@
One notable prior work which applies context-dependent modulation to feedforward networks is context-dependent gating (XdG) (Masse et al., 2018), which implements contextual inputs by concatenation of context signals and gating task-specific neurons. The former could be comparable with our single compartment (basal-only) ablation as it employs a single zone of integration. The main difference with our basal-only ablation is that XdG a) is not informed by intermediate sensory representations, and b) as most prior works, assumes task labels are given and that context always agrees with sensory input. Therefore, it is unknown if XdG, implemented as-is, can deal with cases where context should be overridden by sensory input, as in the unambig.Contra case.
@@
In our experiments, the advantage of apical dendrites was revealed in our multi-scenario contextual integration task. In order to compare our results with previous work, we examined the effect of our integration mechanism in a standard multi-task setting (with task-ids provided to the network, 4). Under these conditions we found that all variants perform similarly well. This can be explained by the setting being simpler, whereas the distinct apical compartment shows its advantage under more varied and challenging types of contextual integration that require inference of ambiguity (no task ids provided).
While not explicitly modular, our framework could be incorporated into the Global Workspace Theory (see (Goyal et al., 2022) for a comparison of recent implementations), to provide a candidate mechanism for resolving inconsistency between modules under ambiguity. It is important to note that our architecture embodying the apical prior presupposes there is asymmetry for how contextual (top-down) and perceptual (bottom-up) representations are combined under ambiguity, which is included in some (e.g. Mittal et al. (2020)) but not other Goyal et al. (2022) implementations of Global Workspace Theory. In general, modular representations can be learned in a distributed and scalable manner, with interactions between local populations (or modules). As highlighted in Goyal et al. (2022), this frees up the capacity of a learning system for more tasks, as it can more efficiently identify task-relevant information and reuse what it has learned. The "apical prior" here implies that these interactions have an asymmetry: they can be contextual signals (in our case via top-down projections to apical compartments), or feed-forward/bottom-up signals. Here we find that this asymmetry provides flexibility to solve perceptual tasks under diverse ambiguity and contextual scenarios.
3.2 What is necessary for robust contextual integration?
@@
Thus far, we have shown that in our contextual integration task, the apical prior provides a specific advantage compared to single-compartment integration when sensory or contextual signals are unreliable. Here, we extend these findings to make broader predictions for what specific components/computations might be necessary for robust contextual integration, and how such a mechanism could plausibly be implemented in the cortex.
@@
Our results suggest that it is sub-optimal to sum all activity together in a single (basal) compartment, if sensory information and context can be contradictory, as would be the case for general cognitive systems. This highlights a key difference between single and two-zone integration of information. As such, an interesting experiment would be to compare the single vs two-zone integration of bottom-up representations and top-down contextual representations in in-vivo experiments or simulations of neurons which more faithfully respect the biophysical constraints of pyramidal neurons in the neocortex, to observe whether our results hold in a more biologically plausible setting, or if it is just a feature of ANNs and/or how they were trained here.
@@
Furthermore, based on our results, we can infer that the apical compartment encodes some information about the ambiguity of the sensory input, and the relevance of context for a given sensory input. An interesting experiment could involve silencing or deactivating apical dendrites in more realistic neurons (biophysical simulations or in-vivo experiments), to assess the effect on accurate perception for our proposed contextual integration tasks.
@@
Our analysis showed that mutual information between apical and basal activation is positive for the ambig.Match scenario. On the other hand, mutual information is close to 0 for the unambig.Irrel and unambig.Contra scenario (Appendix, Table 3). This is not surprising given we explicitly trained under the assumption that local representations, given matching unambiguous contextual information, should collapse to the target representation. Moreover, when sensory ambiguity is low, local representations should be invariant to contextual ambiguity, and mismatching contextual information. This is an important feature of the model supporting robust contextual integration given that most available contextual signals in rich, multi-modal real-world sensory experiences are distractors for any one particular task, and should be discarded.
@@
Following this, we predict that a model cannot learn the relevance of context without encountering and recognizing cases with irrelevant context during training (negative samples). Extending this idea to unsupervised learning, it is interesting to think about how the model could identify contextual relevance. In an unsupervised setting this could arise from an information-maximization learning rule like InfoNCE, which uses a contrastive loss with positive and negative pairs defined via augmentation of the input (positive, relevant) against other inputs (negative, irrelevant). Unfortunately, the creation of positive and negative pairs relies heavily on modality-specific knowledge to construct useful data augmentations (e.g. a rotation applied on an image). A more general approach will be necessary to create such contrastive pairs for many modalities in parallel.
@@
An important feature of our model is that it could learn a mapping of top-down signals onto the apical compartment separately from the contextual representations. This is compatible with regions high in the cortical hierarchy, such as the hippocampus (via the entorhinal cortex), with independent mechanisms for learning a) how they represent contexts, and b) how those contextual representations modulate the sensory representations of lower regions (Maren et al., 2013). Because we assume the stability of the sensory and contextual representation prior to training contextual integration, we cannot make any conclusions from the present study about how these two representations could be simultaneously co-learned. However, we speculate that in the neocortex, there should be some degree of stability in the sensory and contextual representations before learning to perform contextual integration.
This could be compatible with the distributed learning of modules (cortical regions) for sensory and contextual representations initially via self-supervised learning, prior to learning the contextual interactions between the modules (carried by cortico-cortical white matter projections).
@@ 3.3 Limitations
While our model unifies the function of contextual integration with key biophysical properties of dendrites, some of our modeling assumptions and simplifications are worth discussing with respect to their empirical support. Firstly, unlike other computational frameworks such as George et al. (2020) which seek to incorporate the rich diversity and laminar architecture of the neocortex, our model is designed to implement some prominent features of cortical pyramidal neurons without focusing on a specific cortical layer, such as layer 2/3 or layer 5 pyramidal neurons. Specifically, we focus on the functional properties of two-zone dendritic integration. In the cortical circuit, these mechanisms appear to be conserved across L2/3, L5, L6 pyramidal neurons (Larkum et al., 1999; Waters et al., 2003; Larkum et al., 2007; Ledergerber and Larkum, 2010). Here, we assume that contextual signals are provided by neurons in a higher-order region, which would likely correspond to outputs of layer 5 pyramidal neurons in that region, and integrated in the apical dendrites of putative layer 2/3, 4 or 5 pyramidal neurons (Bastos et al., 2012; Schuman et al., 2021). It is worth noting that top-down projections may also innervate the basal compartment of layer 2/3 neurons, and the lower regions of L5 pyramidal neuron basal dendrites (Harris et al., 2019). Despite these missing anatomical features, the framework developed here could serve to study the functional role of laminar architecture in the neocortex: Our study relies largely on the functional mechanism of two-zone integration in pyramidal neurons, which appears to be consistent across pyramidal neuron types, despite potential differences in the sources of the inputs they receive.
Furthermore, our model only implements top-down modulation at a single level in the hierarchy. An obvious next step would be to develop a hierarchical generalization of the approach. While it is unknown how much extra improvement this would provide, evidence of top-down neuron masking (a less expressive form of modulation) at every layer has been shown to improve out-of-distribution robustness (Liu et al., 2023).
One potential criticism is that we train here using gradient descent, for which concerns have been raised as to its biologically plausibility (Lillicrap et al., 2020). The aim of this study is not to find biological learning rules, but rather to understand contextual integration. We use gradient descent as a "best in class" general learning approach to find good solutions to our task, which allows us to analyze the learned representations and the respective contributions of apical and basal compartments.
Our framework lacks the implementation of biological plasticity rules. We leave to future work the exploration of solutions obtained via more biologically plausible plasticity rules such as Lee et al. (2015); Meulemans et al. (2022), compared to those we obtained via backprop. Indeed the analysis above suggests that information-maximization plasticity rules between basal and apical compartments could be a promising direction.
@@ 3.4 Outlook
Often in human brains, inputs from different modalities are processed simultaneously. As a future direction, we propose that our model could be extended to hierarchies of moduels of sensory modalities interacting with each other. Deep learning models that implement multi-modal interactions are an active area of research (Radford et al., 2021; Alayrac et al., 2022; Mustafa et al., 2022). While they excel at multi-modal generation conditioned on one modality, they struggle with conflicting information or contradictions in mixed-modality data (Chen et al., 2024).
As state of the art deep learning models increase exponentially in size, many recent works focus on scalable, parameter-efficient fine-tuning for target tasks, a prominent example being low-rank adapters (LoRAs) Hu et al. (2021); Chen et al. (2024). While LoRAs use resources efficiently, they lack a contextual integration component, and as such, they could be complementary to the "apical prior" architecture studied here. A future direction could combine LoRAs with the apical prior for efficient contextual integration tasks at scale. For example, given that in our task, only a small subset of associations represent matching context, we expect that the weight matrices mapping contextual representations onto apical dendrites could be effectively replaced by LoRAs. This has better scaling properties due to using a smaller fraction of total network capacity.
Initially, stable and useful representations of context could be achieved with self-supervised learning. We speculate that the brain might employ both the apical prior and sparse weights to learn to handle the large pool of available sources of context efficiently. Toward Global Workspace Theory, we argue that a) there must be a mechanism for resolving ambiguity or discrepancy between modules before the step in which modules compete to write to the shared global workspace (as in Goyal et al. (2022)), and b) implementation of the "apical prior" satisfies this requirement. Thus, our present study provides an important step in understanding the process by which brain-wide networks of many modalities can coordinate and integrate information.